ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 85, 1994

Chatzidakis, Z. M. A projective profinite group whose smallest embedding cover is not projective

1

Chatzidakis, Z. M. Some remarks on profinite HNN extensions  11

van den Dries, L. and C. Miller On the real exponential field with restricted analytic functions

19

Efrat, I. and D. Haran On Galois groups over pythagorean and semi-real closed fields

57

Frey, G. Curves with infinitely many points of fixed degree  79

Fried, M. D. and H. Völklein The absolute Galois group of a Hilbertian PRC field

85

Fried, M. D., D. Haran and M. Jarden Effective counting of the points of definable sets over finite fields

103

Geyer, W.-D. On the irreducibility of sums of rational functions with separated variables

135

Haran, D. and L. Lauwers Galois stratification over e-fold ordered Frobenius fields

169

Haran, D. See I. Efrat

Haran, D. See M. D. Fried

Hrushovski, E. A non-PAC field whose maximal purely inseparable extension is PAC

199

Hrushovski, E. and A. Pillay Groups definable in local fields and pseudo-finite fields

203

Jarden, M. The inverse Galois problem over formal power series fields 263

Jarden, M. See M. D. Fried

Kuhlmann, F.-V. Quantifier elimination for henselian fields relative to additive and multiplicative congruences 277

Lubotzky, A. and A. Shalev On some Λ-analytic pro-p groups  307

Lauwers, L. See D. Haran Miller, C. See L. van den Dries Pillay, A. See E. Hrushovski

Schmid, J. On totally positive units of real holomorphy rings  339

Shalev, A. See A. Lubotzky Shelah, S. Vive la différence II. The Ax–Kochen isomorphism

theorem

351

Sonn, J. Brauer groups, embedding problems, and nilpotent groups as Galois groups

391

Völklein, H. Moduli spaces for covers of the Riemann sphere  407

Völklein, H. See M. D. Fried